OpenStudy (itsmichelle29):

Which equation does the graph of the systems of equations solve? two linear functions intersecting at 2, 2 −one halfx + 3 = 3x − 4 −one halfx − 3 = −3x + 4 one halfx + 3 = 3x + 4 one halfx + 3 = −3x − 4

OpenStudy (itsmichelle29):

i medal and fan

OpenStudy (welshfella):

see which of the options fit (2,2) eg first one -1/2 x+ 3 = -1/2* 2 + 3 = ? 3x - 2 = 3(2) - 4 = ? are these 2 equal?

OpenStudy (welshfella):

* typo thats 3x - 4 on 4th line

OpenStudy (itsmichelle29):

NO ?? i think

OpenStudy (campbell_st):

the point (2, 2) needs to be on 2 lines so just substitute and check look at the last choice 1/2(2) + 3 = -3(2) - 4 so this becomes 1 + 3 = -6 -4 or 4 = -10 the solution makes no sense so this isn't the answer start at the top and repeat the process

OpenStudy (welshfella):

what do you get (-1/2)*2 + 3 ?

OpenStudy (welshfella):

= -1 + 3 = ?

OpenStudy (itsmichelle29):

2 lol okay so

OpenStudy (welshfella):

and 3(2) - 4 = ?

OpenStudy (itsmichelle29):

2

OpenStudy (welshfella):

so?

OpenStudy (welshfella):

- you have your answer

OpenStudy (itsmichelle29):

the first one right

OpenStudy (welshfella):

yes

OpenStudy (itsmichelle29):

okay thanks

OpenStudy (welshfella):

- the left side = right side so thats the correct option

OpenStudy (welshfella):

yw